# EMAGE: Towards Unified Holistic Co-Speech Gesture Generation via Expressive Masked Audio Gesture Modeling

### v4

@@ 4.1 Compositional Discrete Face and Body Prior
We model full-body gestures in the separated quantized latent space (c.f. Figure 4) for several reasons. Similar to [64], the body and hands improves the diversity of the results. However, we additionally separate the face and lower body due to their differing correlations with audio. i.e.,, using a single VQ-VAE to encode both upper and lower body may lead the model to overlook gestures that occur less frequently - regardless of their relation to the audio. Specifically, a single model might focus only on recovering lower body motion, neglecting elbow movements in a speaker who is constantly walking during the conversation.
The separated quantized latent space $Q=\{\mathbf{q}_{\text{f}},\mathbf{q}_{u},\mathbf{q}_{\text{h}},\mathbf{q}_{\text{l}}\}$ for face $\mathbf{g}_{\text{f}}\in\mathbb{R}^{T\times 106}$, Upper body $\mathbf{g}_{\text{u}}\in\mathbb{R}^{T\times 78}$, Hands $\mathbf{g}_{\text{h}}\in\mathbb{R}^{T\times 180}$, and Lower body $\mathbf{g}_{\text{l}}\in\mathbb{R}^{T\times(54+4)}$ are from four VQ-VAEs. The VQ-VAE for each is optimized by jointly optimizing the following loss terms,
@@
where $\mathbf{z}$ is the encoded $\mathbf{g}$ with a temporal window size $w=1$. $\mathcal{L}_{\text{rec}}$, $\mathcal{L}_{\text{vel}}$, and $\mathcal{L}_{\text{acc}}$ are Geodesic [57] and L1 loss, sg is a stop gradient operation, we set the weight of the commitment loss [58] (the last term) to 1 in this paper.
@@ 4.2 Content and Rhythm Self-Attention
Given the speech audio $\mathbf{s}$, inspired by [5], we employ onset $\mathbf{o}$ and amplitude $\mathbf{a}$ as explicit audio rhythm, alongside the pretrained embeddings [9] $\mathbf{e}$ from transcripts as content. Different from previous approaches, which typically add the rhythm and content features, we leverage self-attention to merge these features adaptively. This approach is driven by the observation that for specific frames, the gestures are more related to content (semantic-aware) or rhythm (beat-aware). The rhythm and content features are first encoded into time-aligned features $\mathbf{r}_{1:T}$ and $\mathbf{c}_{1:T}$, using a Temporal Convolutional Network (TCN) and linear mapping, respectively. For each time step $t\in\{1,\dots,T\}$, we merge the rhythm and content features by:
@@ 4.3 Masked Audio-Conditioned Gesture Modeling
We propose a Masked Audio Gesture Transformer to leverage different training paths (cf. Fig. 5 for the motivation behind the architecture designs). Given temporal and spatially masked gestures $\overline{\mathbf{g}}\in\mathbb{R}^{T\times 337}$, we first replace the masked tokens with learned masked embeddings $e_{\text{mask}}\in\mathbb{R}^{256}$, as the value zero still represents specific motion content, e.g., a T-pose. We linearly increase the ratio of masked joints and frames from 0 to 95% according to the training epochs.
A spatial convolutional $\mathcal{SC}$ encoder initially summarizes the spatial information, and compress the spatial feature into $\mathbb{R}^{T\times 256}$. We then employ a temporal self-attention (without feed-forward) $\mathcal{TSA}$ to refine the summarized spatial features. A straightforward Teacher-Force temporal cross-attention Transformer decoder $\mathcal{TCAT}$ is adopted for gesture reconstruction. It takes in the $\overline{\mathbf{g}}$ and $e_{\text{mask}}$, outputting the reconstructed latent $\hat{\mathbf{l}}_{\text{mg2g}}$,
$\mathbf{h}=\mathcal{TSA}(\mathcal{SC}(\overline{\mathbf{g}}+\mathbf{p}_{s})+\mathbf{p}_{t}),$ (6)
where $\mathbf{h}\in\mathbb{R}^{T\times 512}$ represents encoded body hints, $\mathbf{p}_{s}$ is a learned spatial position embedding, and $\mathbf{p}_{t}$ is the sum of a learned speaker embedding and PPE [20].
@@
Then, we minimize the L1 distance in the latent space,
@@
We optimize both the latent code class classification cross-entropy loss $\mathcal{L}_{\text{a2g-rec}}$ and the latent reconstruction loss $\mathcal{L}_{\text{a2g-cls}}$ to encourage the results’ diversity.
@@ 4.4 Face and Translation Decoding
AMASS[45] SMPLX FLAME 42.0 ± 3.60 -
TalkShow[64] SMPLX SMPLX 14.4 ± 2.19 26.1 ± 6.42
BEAT2 (Ours) SMPLX FLAME 43.6 ± 3.38 35.7 ± 5.91
@@ 5.1 Dataset Quality Evaluation
We compare our dataset with the state-of-the-art pseudo ground truth (PGT) dataset, Talkshow [64], for both face and body. Additionally, we compare it with the AMASS [45] dataset for the body and VOCA [14] for the face as references. Due to the varying lengths of sequences in each dataset, we sample 100 comparison pairs, each with an equal duration ranging from 2 to 4 seconds. Each user evaluates a random set of 40 pairs during a 10-minute session by selecting the sequence they consider to have the best captured quality. In total, 60 users were invited to participate. It is important to note that users are instructed to compare only the upper body results, as Talkshow contains only the upper body. The results are shown in Table 3.
@@ 5.2 Model Ability Evaluation
We report Table 4 on BEATv2 and ablation studies on BEATv1.3, an earlier release version of BEATv2. Compared to BEATv1.3, BEATv2 refines the facial data by animators and filters the hand data by human annotators.
@@
Metrics
We adopt FGD[65] to evaluate the realism of the body gestures. Then, we measure Diversity[33] by calculating the average L1 distance between multiple body gesture clips, and use BC [34] to assess speech-motion synchrony. For the face, we calculate the vertex MSE[60] to measure the positional distance and the vertex L1 difference LVD [64] between the GT and generated facial vertices.
Compared Methods
We first compare our method with representative state-of-the-art methods in body gesture generation [25, 65, 41, 36, 39, 61] and talking head generation [20, 60] by reproducing their methods for body and face, respectively. In addition, we reproduce two previous state-of-the-art holistic pipelines, Habibie et al. [28] and Talkshow [64], whose original implementations are limited to the upper body. We adds a lower-body decoder to Habibie’s method and a lower-body VQ-VAE to Talkshow’s method.
@@ 5.2.1 Quantitative and Qualitative Analysis
As shown in Table 4, with a four-frame seed pose, our method outperforms previous SOTA algorithms. For qualitative results, we first present subjective evaluations in Figures 6 and G. Furthermore, we conduct a perceptual user study. Maintaining the same setup of 60 users, each user evaluates 40 pairs of 10-second results. Users decide which result is more believable as the win rate, with results shown in Table 5.
Figure 7: Comparison of Generated Facial Movements. Compared with previous state-of-the-art talking face generation methods faceformer[20] and codetalker[60], holistic gestures generation methods habibie et al.[28] and Talkshow [64], notes that in BEAT2, codetalker higher MSE (lower is better) but better subjective results as discrete face prior. but due to our method present more accurate lip motions by leveraging the advantage of discrete face prior and masked body hints.
@@ 5.2.2 Ablation Analysis
Performance of Baseline
As shown in Table 6, we start with a teacher-force transformer-based baseline. This baseline, inspired by FaceFormer [20], adopts a 1-layer cross-attention transformer decoder and replaces the audio features from Wav2Vec2 [6] with our custom TCN [7] and trainable word embeddings [9].
Effect of Multiple VQ-VAEs
Simply applying one VQ-VAE [58, 27] for full-body movements, including the face, decreases performance in facial movements. This is because the VQ-VAE is trained to minimize the average loss for the full body, and some speakers’ most frequent movements are not related to the audio. Implementing separated VQ-VAEs allows the model to better leverage the advantages of discrete priors.
Effect of Content Rhythm Self-Attention
Adaptive fusion of rhythm and content features shows improvement in both FGD and Alignment. It selectively focuses the current motion more on rhythm or content features based on the training data’s distribution. Moreover, we observed more semantic-aware results when applying CRA.
Effect of Masked Body Gesture Hints
The improvement across all objective metrics demonstrates that our model effectively leverages spatial-temporal gesture priors, reducing the likelihood of incorrect gesture sampling. Furthermore, masked gesture modeling is key to enabling the network to accept customized inputs in the middle frames.
Table 6: Abliation Analysis on BEATv1.3
@@
We demonstrate that EMAGE can effectively combine multiple non-holistic datasets for training by jointly training with the retargeted Trinity [22] and AMASS [45], accepting only the upper body and audio pairs from Trinity and only body and hands from AMASS. Separate VQ-VAEs were trained for BEAT2, Trinity, and AMASS, with separate MLP heads implemented for codebook classification. The results in Table 7 show that incorporating data improves performance on the BEATv1.3 test set.
@@ 6 Conclusion
In this work, we present EMAGE, a framework designed to accept partial gestures as input for completing audio-synchronized holistic gestures. It demonstrates that leveraging masked gesture reconstruction can significantly enhance the performance of audio-conditioned gesture generation. Furthermore, the design of EMAGE enables the training on multiple datasets, further improving performance. Alongside EMAGE, we release BEAT2, the largest mesh dataset adhering to academic standards in both face and body formats. We hope that BEAT2 will contribute to knowledge and model sharing across various subareas.
7 Conflict of Interest Disclosure
While You Zhou, Xuefei Zhe, Naoya Iwamoto, and Bo Zheng are employees of Huawei Tokyo Research Center, this work was done on their own time with the approval of their employer.
MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon. MJB has financial interests in Amazon and Meshcapade GmbH. While MJB is a co-founder and Chief Scientist at Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.

### v5

@@ 4.1 Compositional Discrete Face and Body Prior
We model full-body gestures in the separated quantized latent space (Figure 4) for several reasons. Similar to [65], the body and hands improves the diversity of the results. However, we additionally separate the face and lower body due to their differing correlations with audio, i.e., using a single VQ-VAE to encode both the upper and lower body may lead the model to overlook gestures that occur less frequently - regardless of their relation to the audio. Specifically, a single model might focus only on recovering lower body motion, neglecting elbow movements in a speaker who is constantly walking during the conversation.
The separated quantized latent space $Q=\{\mathbf{q}_{\text{f}},\mathbf{q}_{u},\mathbf{q}_{\text{h}},\mathbf{q}_{\text{l}}\}$ for the Face $\mathbf{g}_{\text{f}}\in\mathbb{R}^{T\times 106}$, Upper body $\mathbf{g}_{\text{u}}\in\mathbb{R}^{T\times 78}$, Hands $\mathbf{g}_{\text{h}}\in\mathbb{R}^{T\times 180}$, and Lower body $\mathbf{g}_{\text{l}}\in\mathbb{R}^{T\times(54+4)}$ are from four VQ-VAEs. The VQ-VAE for each is optimized by jointly optimizing the following loss terms,
@@
where $\mathbf{z}$ is the encoded $\mathbf{g}$ with a temporal window size $w=1$. $\mathcal{L}_{\text{rec}}$, $\mathcal{L}_{\text{vel}}$, and $\mathcal{L}_{\text{acc}}$ are Geodesic [58] and L1 losses, sg is a stop gradient operation, we set the weight of the commitment loss [59] (the last term) to 1 in this paper.
@@ 4.2 Content and Rhythm Self-Attention
Given the speech audio, $\mathbf{s}$, inspired by [4], we employ onset $\mathbf{o}$ and amplitude $\mathbf{a}$ as explicit audio rhythm, alongside the pretrained embeddings [9] $\mathbf{e}$ from transcripts as content. Different from previous approaches, which typically add the rhythm and content features, we leverage self-attention to merge these features adaptively. This approach is driven by the observation that for specific frames, the gestures are more related to content (semantic-aware) or rhythm (beat-aware). The rhythm and content features are first encoded into time-aligned features $\mathbf{r}_{1:T}$ and $\mathbf{c}_{1:T}$, using a Temporal Convolutional Network (TCN) and linear mapping, respectively. For each time step $t\in\{1,\dots,T\}$, we merge the rhythm and content features by:
@@ 4.3 Masked Audio-Conditioned Gesture Modeling
We propose a Masked Audio Gesture Transformer to leverage different training paths (see Fig. 5 for the motivation behind the architecture designs). Given temporal and spatially masked gestures $\overline{\mathbf{g}}\in\mathbb{R}^{T\times 337}$, we first replace the masked tokens with learned masked embeddings $e_{\text{mask}}\in\mathbb{R}^{256}$, as the value zero still represents specific motion content, e.g., a T-pose. We linearly increase the ratio of masked joints and frames from 0 to 95% according to the training epochs.
A spatial convolutional $\mathcal{SC}$ encoder initially summarizes the spatial information, and compress the spatial feature into $\mathbb{R}^{T\times 256}$. We then employ a temporal self-attention (without feed-forward) $\mathcal{TSA}$ to refine the summarized spatial features
$\mathbf{h}=\mathcal{TSA}(\mathcal{SC}(\overline{\mathbf{g}})+\mathbf{p}_{t}),$ (6)
where $\mathbf{h}\in\mathbb{R}^{T\times 512}$ represents encoded body hints, and $\mathbf{p}_{t}$ is the sum of a learned speaker embedding and PPE [19]. Then, a straightforward temporal cross-attention Transformer decoder $\mathcal{TCAT}$ is adopted for the reconstructed latent $\hat{\mathbf{q}}_{\text{mg2g}}$,
@@
We minimize the L1 distance in the latent space,
@@
We optimize both the latent code class classification cross-entropy loss $\mathcal{L}_{\text{a2g-rec}}$ and the latent reconstruction loss $\mathcal{L}_{\text{a2g-cls}}$ to encourage diverse results.
@@ 4.4 Face and Translation Decoding
AMASS[46] SMPL-X FLAME 42.0 ± 3.60 -
TalkSHOW[65] SMPL-X SMPL-X 14.4 ± 2.19 26.1 ± 6.42
BEAT2 (Ours) SMPL-X FLAME 43.6 ± 3.38 35.7 ± 5.91
@@ 5.1 Dataset Quality Evaluation
We compare our dataset with the state-of-the-art pseudo ground truth (PGT) dataset, TalkSHOW [65], for both face and body. Additionally, we compare it with the AMASS [46] dataset for the body and VOCA [14] for the face as references. Due to the varying lengths of sequences in each dataset, we sample 100 comparison pairs, each with an equal duration ranging from 2 to 4 seconds. In a perceptual study, each participant evaluates a random set of 40 pairs during a 10-minute session by selecting the sequence they consider to have the best captured quality. In total, 60 participants were invited. It is important to note that participants are instructed to compare only the upper body results, as TalkSHOW contains only the upper body. The results are shown in Table 3.
@@ 5.2 Model Ability Evaluation
We report results using BEATv1.3 and BEATv2. The latter has facial data refined by animators and hand data filtered by annotators.
@@
Metrics. We adopt FGD[67] to evaluate the realism of the body gestures. Then, we measure Diversity[33] by calculating the average L1 distance between multiple body gesture clips, and use BC [34] to assess speech-motion synchrony. For the face, we calculate the vertex MSE[61] to measure the positional distance and the vertex L1 difference LVD [65] between the GT and generated facial vertices.
Compared Methods. We first compare our method with representative state-of-the-art methods in body gesture generation [25, 67, 40, 38, 39, 62] and talking head generation [19, 61] by reproducing their methods for body and face, respectively. In addition, we reproduce two previous state-of-the-art holistic pipelines, Habibie et al. [28] and TalkSHOW [65], whose original implementations are limited to the upper body. We add a lower-body decoder to Habibie et al. and a lower-body VQ-VAE to TalkSHOW.
@@ 5.2.1 Quantitative and Qualitative Analysis
As shown in Table 4, with a four-frame seed pose, our method outperforms previous SOTA algorithms. For qualitative results, see Figures 6 and 7. Furthermore, we conduct a perceptual study. Maintaining the same setup of 60 participants, each participant evaluates 40 pairs of 10-second results do decide which is most believable; this gives the win rate shown in Table 5.
Figure 7: Comparison of Generated Facial Movements. Compared with previous state-of-the-art talking face generation methods FaceFormer [19] and CodeTalker [61] as well as holistic gesture generation methods Habibie et al. [28] and TalkSHOW [65]. Note that CodeTalker has higher MSE than EMAGE on BEATv2 (Table 4, lower is better) but is subjectively realistic. EMAGE gets good lip motions by leveraging both the face model and masked body hints.
@@ 5.2.2 Ablation Analysis
Performance of Baseline. As shown in Table 6, we start with a teacher-force transformer-based baseline. This baseline, inspired by FaceFormer [19], adopts a 1-layer cross-attention transformer decoder and replaces the audio features from Wav2Vec2 [6] with our custom TCN [7] and trainable word embeddings [9].
Effect of Multiple VQ-VAEs. Simply applying one VQ-VAE [59, 27] for full-body movements, including the face, decreases performance in facial movements. This is because the VQ-VAE is trained to minimize the average loss for the full body, and some speakers’ most frequent movements are not related to the audio. Implementing separated VQ-VAEs allows the model to better leverage the advantages of discrete priors.
Effect of Content Rhythm Self-Attention. Adaptive fusion of rhythm and content features shows improvement in both FGD and Alignment. It selectively focuses the current motion more on rhythm or content features based on the training data’s distribution. Moreover, we observed more semantic-aware results when applying CRA.
Effect of Masked Body Gesture Hints. The improvement across all objective metrics demonstrates that our model effectively leverages spatial-temporal gesture priors, reducing the likelihood of incorrect gesture sampling. Furthermore, masked gesture modeling is key to enabling the network to accept predefined gestures in specific frames.
@@
We demonstrate that EMAGE can effectively combine multiple non-holistic datasets for training by jointly training with the Trinity [22] and AMASS [46] datasets, by using only the upper body and audio pairs from Trinity and only body and hands from AMASS. Separate VQ-VAEs are trained for BEAT2, Trinity, and AMASS, with separate MLP heads implemented for codebook classification. The results in Table 7 show that incorporating data improves performance on the BEATv1.3 test set.
@@ 6 Conclusion
In this work, we present EMAGE, a framework designed to accept partial gestures as input for completing audio-synchronized holistic gestures. It demonstrates that leveraging masked gesture reconstruction can significantly enhance the performance of audio-conditioned gesture generation. Furthermore, the design of EMAGE enables the training on multiple datasets, further improving performance. Alongside EMAGE, we release BEAT2, the largest multi-modal gesture dataset consistent with SMPL-X and FLAME. We hope that BEAT2 will contribute to knowledge and model sharing across various subareas.
Disclosures: While You Zhou, Xuefei Zhe, Naoya Iwamoto, and Bo Zheng are employees of Huawei Tokyo Research Center, this work was done on their own time with the approval of their employer. MJB CoI: https://files.is.tue.mpg.de/black/CoI_CVPR_2024.txt